OpenStudy (anonymous):

what is the limit of (1/x - 1/2) / x-2 as x approaches 2?

OpenStudy (anonymous):

\[\lim \frac{ \frac{ 1 }{ x } - \frac{ 1 }{ 2 } }{ x-2 }\]

OpenStudy (anonymous):

as x approaches 2

OpenStudy (anonymous):

first simplify the equation or get rid off the fraction in the numerator by multiplying both denominator and numerator by the common denominator of the numerator..

OpenStudy (anonymous):